OpenStudy (anonymous):

Which statement is true? A. –14 + 21 = –24 + 17 B. 16 + (–6) > 24 + (–13) C. –14 + (–10) > 2 + (–28) D. –10 + (–8) > –12 + (–4)

OpenStudy (31356):

What do you think? o.0

OpenStudy (anonymous):

im not sure :T

OpenStudy (anonymous):

what do you think?

OpenStudy (31356):

It's not A

OpenStudy (31356):

Insert it in the calculator and observe :D

OpenStudy (anonymous):

yea that's what i'm doin

OpenStudy (31356):

Thx for the medal :D

OpenStudy (31356):

Not B either

OpenStudy (anonymous):

i think its D

OpenStudy (doc.brown):

Sorry, you have to do each one, best to write it down\[A\]\[–14 + 21 = –24 + 17\]\[7=-7\]Not true, next?

OpenStudy (31356):

It's C :D

OpenStudy (anonymous):

ohhh

OpenStudy (31356):

-14+9-10)= -24 2+(-28) = -26 -24 is greater than -26 :D Remember it is opposite in the negatives

OpenStudy (31356):

Medal? If helped........

OpenStudy (anonymous):

im horrable at math :T cant really help

OpenStudy (31356):

Lol, you can do it :D Practice, practice makes perfect

OpenStudy (anonymous):

same thing my mom says lol

OpenStudy (anonymous):

What is the value of the variable that makes the statement true? c + (–8) + (–5) = 0 A. –13 B. –3 C. 3 D. 13

OpenStudy (31356):

Hmm....

OpenStudy (31356):

Not a

OpenStudy (31356):

Not b

OpenStudy (31356):

It's D :D

OpenStudy (anonymous):

for yours?

OpenStudy (31356):

Do calculator check D:

OpenStudy (31356):

Thx for the medal :D

OpenStudy (anonymous):

yw :)

OpenStudy (31356):

Triple smileys :D:D:D lol

OpenStudy (anonymous):

What is the value of |x| + y when x = –3 and y = 15? A. –18 B. –12 C. 12 D. 18

OpenStudy (31356):

Hmm.....

OpenStudy (anonymous):

i believe it's c wanna be sure

OpenStudy (31356):

Yea it's, I checked it on my calculator :D

OpenStudy (anonymous):

oki :D:D

OpenStudy (31356):

:D:D:D

OpenStudy (anonymous):

:D Mastered

OpenStudy (31356):

Good! 100% right? :D

OpenStudy (anonymous):

Yea :D